# QG from SymQRG: AdS3/CFT2 Correspondence as Topological Symmetry-Preserving Quantum RG Flow

### v1

@@ 6.4.2 Chaos from (topological) Order
in the thermal state with inverse temperature $\beta$. It is crucial that the four operators are in Lorentzian out-of-time ordering. In fact, if the four operators are all in Euclidean signature, the ordering of the operators does not matter at all, and we can even use this single-valuedness as one of the bootstrap equations. The Lorentzian correlators are all analytical continuation of Euclidean ones, and each choice of the ordering in Lorentzian signature correlators correspond to a specific $i\epsilon$ prescription [254, 255, *Cornalba:2006xm, *Cornalba:2007zb, *Cornalba:2007fs, 251] [^44] Study of Lorentzian CFT in the analytical conformal bootstrap program in recent years have lead to fruitful new results, see [259].. More specifically, the CFT four point functions in Euclidean signature can be expanded in terms of the conformal blocks as:
@@
In Euclidean signature, the cross ratios $z$ and $\bar{z}$ are complex conjugates of each other, shown as the black dots. In Lorentzian signature, $z$ and $\bar{z}$ are independent, and also, when we carefully follow the $i\epsilon$ prescription in Lorentzian signature, we see that the cross ratio $z$ in the OTOC crosses the branch cut at $z=1$, which is usually said to “go to the second sheet”, and then land on the blue dot on the second sheet. Notice that $\bar{z}$ however does not cross any branch cuts, so there is an asymmetry between $z$ and $\bar{z}$, which is important to get a different full correlator in Lorentzian signature. The branch cut in the cross ratio space for the conformal blocks also has an interpretation in real space[260, 261], as coming from the $W$ operator entering the lightcone of the $V$ operator, and the difference in the choice of branch cuts is what gives the non-vanishing commutator for timelike separated operators, and is intrinsic to causality and Lorentzian signature. The “Lyapunov regime” common in the study of chaos further corresponds to taking $z,\bar{z}\to 0$, and this kinematical regime is usually called the CFT Regge limit[255, *Cornalba:2006xm, *Cornalba:2007zb, *Cornalba:2007fs, 262]. Assuming Virasoro identity block dominance and using the explicit expression for the identity block in the heavy-heavy-light-light limit, we can get the maximal Lyapunov exponent behavior expected for 2D holographic CFTs[263, 264].
In fact, there is an exact answer for the OTOC, without using any assumptions, using the fact that the winding around branch point $z=1$ is picking up a monodromy of the conformal blocks, and there is a monodromy matrix $M$ implementing it exactly[52, *Moore:1989vd], such that
@@
The fact that we can write the OTOC in terms of the topological data is certainly not just a coincidence. It is the fact that the monodromy is just a combination of braid moves, as shown in Fig. 29464646Here, we are employing a reversed logic compared to [52, 60]: Suppose we already possess the Moore-Seiberg data of the CFT. We can use this data to determine the conformal blocks on the second sheet. Notably, this data was originally derived by first analyzing the monodromies of the conformal blocks.. From the bulk-boundary correspondence[94], or unfolding the sandwich construction, we can also see the monodromy explicitly as anyons nontrivially linking themselves in the bulk, as shown in [265, *Gu:2016hoy].
@@
This point of view of understanding quantum chaos and OTOC from the exchange algebra was emphasized in the context of 2D CFT in [63, 267, 268]. In the context of holography, this is the dual of the bulk exchange algebra found in [269, *Kiem:1995iy], responsible for the exponential blueshift coming from the Dray and ’t Hooft shockwave effect[271]. As noted in [63], the goal of conformal bootstrap is to ensure that correlation functions remain single-valued in Euclidean signature. In other words, quantum statistics are invisible or “confined” in Euclidean, non-chiral, full CFTs. However, after we solve the bootstrap equations which are topological in nature as we explained, and manage to hide the quantum statistics, these topological data will get “deconfined” in Lorentzian signature, controlling the Lorentzian dynamics, causality(commutator), and information scrambling of the theory, which gives “Chaos from Order”.
@@ 7 Discussions
As discussed above, the bulk dual of the CFT is represented by a sum over quantum geometries, providing an exact description at finite $N$. In the large $N$ limit, certain quantum geometries emerge as dominant saddles in this sum, allowing for a semi-classical, low-energy geometrical description. To better understand this emergence, recent studies have proposed algebraic quantum field theory (AQFT), specifically von Neumann algebras, as a suitable framework, noting the appearance of emergent Type III von Neumann algebra structures in the large $N$ limit[273, *Leutheusser:2021frk].
A natural question arises: how are these two distinct algebraic approaches for field theories and emergence of geometry related, and how are categorical structures, such as the $6j$ symbols, encoded within von Neumann algebras? In 2D CFT, there appears to be a dictionary connecting these approaches. AQFT in 2D is rigorously constructed using “conformal nets”, and its connection to the functorial formalism is discussed in [275]. A key element in this framework is the role of superselection sectors and “subfactor theory” [276, *Longo:1989tt] [^48] We thank Keiichiro Furuya for discussions on this topic and collaboration on related project.. This approach might provide a new understanding of the Type II algebra structures observed in the context of higher-dimensional gravity [278], which plays a role in black holes and cosmology [279, *Chandrasekaran:2022cip, 281]. Applied to 3D gravity, this Type II algebra is analogous to the JT gravity case studied in [282, *Kolchmeyer:2023gwa]. In the current situation, we even achieve its UV completion, since we start from a 2D CFT. We will report further results on these ideas in future work.
@@
In the case of irrational CFTs, an additional layer of infinity appears due to the infinite number of primary fields in |$\ket{\Psi}$. This infinity drives the large $N$ limit of a CFT and enables chaotic behavior, in contrast to rational CFTs. It mirrors the emergence of Type III factors in the large $N$ limit as discussed above [273, *Leutheusser:2021frk]. For the SymTFT description, this implies that an infinite number of anyons are required to describe the bulk theory of gravity. Consequently, for irrational CFTs, we do not have a fusion category, which by definition requires finitely many objects. This fact also limits our Turaev-Viro construction for $U_{q}(SL(2,\mathbb{R}))$, as it lacks the 4-1 move invariance. However, when we consider these three-dimensional manifolds with boundaries, combined with boundary conformal blocks, the construction does make sense and yields a convergent result consistent with the Virasoro/Teichmuller TQFT, at least for the diagonal Liouville CFT, as explicitly shown in [1, 35]. This suggests the existence of a novel mathematical structure, essential for the study of quantum gravity.
@@
Our tensor network formalism depends significantly on the unique properties of 3D gravity, particularly its topological nature. Extending this line of reasoning may provide insight into constructing a non-perturbative bulk dual in higher dimensions. However, since gravitons are dynamical in higher dimensions, we want to explore several possibilities to address this tension. One possibility, as noted in [47], is that the bulk theory may lose its topological character due to the irrational nature of the symmetry. Another possibility is that the bulk topological invariance could be spontaneously broken, even if it initially exhibits a topological structure. Alternatively, it is possible that the bulk theory remains genuinely topological in the non-perturbative regime.
@@
The volume/action of the bulk theory is proposed to be connected to the notion of complexity in the CFT [288, *Stanford:2014jda, *Brown:2015bva]. There is a large volume of work exploring this connection, and it has been shown that perhaps more generally there are an infinite class of bulk quantities that could be understood as complexity[291, *Belin:2022xmt, *Myers:2024vve]. However, precise definitions of complexity on both sides have been tricky to identify. Our discrete formulation of the path-integral could be used to define notions of complexity in field theories that can be directly translated between the bulk and the boundary. For example, since each 6j symbol for the Virasoro lines can be interpreted as the volume of hyperbolic tetrahedra, we can assign definite complexity to individual “gate” corresponding to the 6j symbols that builds up the path-integral. Our $U_{q}(SL(2,\mathbb{R}))$ $6j$ symbols seem to be particularly promising, since Liouville theory has be proposed to capture path integral complexity for 2D CFTs[294, *Caputa:2017yrh].
@@ Acknowledgement
We thank Nathan Banjamin, Christian Ferko, Keiichiro Furuya, Hao Geng, Sarah Harrison, Daniel Jafferis, Hong Liu, Juan Maldacena, Shinsei Ryu, Sahand Seifnashri, Shu-Heng Shao, Herman Verlinde and Frank Verstraete for helpful discussions. We thank Wei Bu, Wan Zhen Chua, Christian Ferko, Hao Geng, Ziming Ji, Chen-Te Ma, Yixu Wang, Zixia Wei and Chen Yang for comments and discussions on the draft. We thank Lin Chen, Gong Cheng, Keiichiro Furuya, Hao Geng, Zheng-Cheng Gu, Bing-Xin Lao for collaboration on related projects. We thank Zhenhao Zhou for plotting the 3D diagrams in this paper. YJ thanks Juan Maldacena for the invitation and hospitality during a visit to the Institute for Advanced Study, where the idea of this paper originated from discussions with people there. YJ and NB acknowledge the support by the U.S Department of Energy ASCR EXPRESS grant, Novel Quantum Algorithms from Fast Classical Transforms, and Northeastern University. LYH acknowledges the support of NSFC (Grant No. 11922502, 11875111). ZL is supported by NSF grant PHY-2110463.
@@ A.1 Turaev-Viro TQFT from Chern-Simons theory
By considering the dual graph corresponding to the network of Wilson lines, and recursively using the relation in Fig. 33, it recovers the path integral with the insertion of $\omega$ loops, which serves only to project every loop back to vanishing flux[298, *2011NJPh...13f5001B]. The resultant network is called the “Chain-Mail” [296]. The corresponding expectation value of closed Wilson loops is denoted as $C(\mathcal{M},D)$. Roberts invariants can be defined as:
@@ A.2 Turaev-Viro TQFT from fusion category
Generically, it is possible to endow a fusion category with more structures. An important structure is the braided structure that exchanges two objects. This structure is encoded in the so-called $R$-matrix, with components $R^{c}_{ab}$ which are phases, that connects $a\otimes b$ to $b\otimes a$. This is illustrated in Fig. 37. Note that there is a distinction between an “over-crossing” and “under-crossing”. The factor that appears in an “under-crossing” is related to the “over-crossing” by a complex conjugate of $R_{ab}^{c}$. The $R$-matrix satisfies some consistency conditions, so that the action of braiding is consistent with the associativity relations from fusion. These consistency conditions are given by the Hexagon relations[60, *Moore:1989vd].
@@ Appendix B Tensor network state sum representation of 2D CFT path integrals
For the computation of the total CFT partition $\mathcal{Z}_{\text{CFT}}$ function, the basic building block is the CFT path integral on the chipped triangles, as shown in Fig. 9525252We will focus on 2D CFT defined on flat background $\delta_{ij}$ for simplicity here. The formalism can be easily generalized to general fixed curved backgrounds $h_{ij}$.. These represent open pairs of pants, in contrast to the more conventional closed pairs of pants commonly found in the CFT literature. The Hilbert spaces in the computation are associated with the edges of the triangles, which are lines between the two endpoints where boundary conditions must be specified.
@@
For diagonal CFTs, the label set of conformal boundary conditions coincides with the set of primary fields (this is referred to as the Cardy case [107] in the literature), and the rescaled BCFT OPE coefficients are directly proportional to the quantum $6j$ symbols, as first discovered by Runkel [103, *Runkel:1999dz], as
@@
where $d_{\alpha_{i}}$ is the quantum dimension associated with $\alpha_{i}$, as introduced above. For theories with a continuous spectrum, such as the Liouville theory, this corresponds to the Plancherel measure [1, 160, *Ponsot:2000mt]. In the diagonal case, the $6j$ symbol enjoys the tetrahedron symmetry, and we associate it with a tetrahedron, which serves as the fundamental building block of the Turaev-Viro TQFT, as reviewed above.
@@ Appendix C Tensor network SymQRG
The orthogonality condition of the (generalized) $6j$ symbols is[36, *Fuchs:2004xi, 105],
@@
The next step is to express the wavefunctions as contractions of local tensors defined on the triangles as in Fig. 45, using singular value decomposition [113, *PhysRevLett.115.180405, 33], as follows:
@@
In the context of 3D gravity, the idea of seeing the emergence of gravity by treating closed OPE coefficients in the pair-of-pants decomposition as random tensors was introduced in [307, *Verlindeunpublished], and later adapted into a tensor network framework in [309]. For example, a simple MPS tensor network, constructed with random closed OPE coefficients in a 2D CFT as tensors, can be interpreted as discretizing the radial direction in the bulk dual in the semi-classical approximation, and approximately reproduces the location of the horizon[307, *Verlindeunpublished, 309]. In our framework, which recovers the full CFT path integral, it is also natural to consider an ensemble average of the structure coefficients for boundary-changing operators.
@@
The ensemble average of the structure coefficients for the boundary-changing operators in the large central charge limit has been studied in [191, *Kusuki:2022wns, 193]. Applied to our formalism, this results in a random tensor network that represents the ensemble-averaged CFT path integral as outlined in [1], and will appear soon in [195]. A crucial distinction in our tensor network construction is worth emphasizing: We used the open structure coefficients of the CFT, rather than the closed ones, to build the CFT tensor network. From our perspective, the open structure coefficients are more directly connected to the emergence of bulk geometry and allow for discretizing both the CFT and AdS spaces arbitrarily, not only along the radial direction or having to come with external operator insertions.

### v2

@@ 6.4.2 Chaos from (topological) Order
in the thermal state with inverse temperature $\beta$. It is crucial that the four operators are in Lorentzian out-of-time ordering. In fact, if the four operators are all in Euclidean signature, the ordering of the operators does not matter at all, and we can even use this single-valuedness as one of the bootstrap equations. The Lorentzian correlators are all analytical continuation of Euclidean ones, and each choice of the ordering in Lorentzian signature correlators correspond to a specific $i\epsilon$ prescription [263, 264, *Cornalba:2006xm, *Cornalba:2007zb, *Cornalba:2007fs, 260] [^47] Study of Lorentzian CFT in the analytical conformal bootstrap program in recent years have lead to fruitful new results, see [268].. More specifically, the CFT four point functions in Euclidean signature can be expanded in terms of the conformal blocks as:
@@
In Euclidean signature, the cross ratios $z$ and $\bar{z}$ are complex conjugates of each other, shown as the black dots. In Lorentzian signature, $z$ and $\bar{z}$ are independent, and also, when we carefully follow the $i\epsilon$ prescription in Lorentzian signature, we see that the cross ratio $z$ in the OTOC crosses the branch cut at $z=1$, which is usually said to “go to the second sheet”, and then land on the blue dot on the second sheet. Notice that $\bar{z}$ however does not cross any branch cuts, so there is an asymmetry between $z$ and $\bar{z}$, which is important to get a different full correlator in Lorentzian signature. The branch cut in the cross ratio space for the conformal blocks also has an interpretation in real space[269, 270], as coming from the $W$ operator entering the lightcone of the $V$ operator, and the difference in the choice of branch cuts is what gives the non-vanishing commutator for timelike separated operators, and is intrinsic to causality and Lorentzian signature. The “Lyapunov regime” common in the study of chaos further corresponds to taking $z,\bar{z}\to 0$, and this kinematical regime is usually called the CFT Regge limit[264, *Cornalba:2006xm, *Cornalba:2007zb, *Cornalba:2007fs, 271]. Assuming Virasoro identity block dominance and using the explicit expression for the identity block in the heavy-heavy-light-light limit, we can get the maximal Lyapunov exponent behavior expected for 2D holographic CFTs[272, 273].
In fact, there is an exact answer for the OTOC, without using any assumptions, using the fact that the winding around branch point $z=1$ is picking up a monodromy of the conformal blocks, and there is a monodromy matrix $M$ implementing it exactly[44, *Moore:1989vd], such that
@@
The fact that we can write the OTOC in terms of the topological data is certainly not just a coincidence. It is the fact that the monodromy is just a combination of braid moves, as shown in Fig. 29494949Here, we are employing a reversed logic compared to [44, 52]: Suppose we already possess the Moore-Seiberg data of the CFT. We can use this data to determine the conformal blocks on the second sheet. Notably, this data was originally derived by first analyzing the monodromies of the conformal blocks.. From the bulk-boundary correspondence[72], or unfolding the sandwich construction, we can also see the monodromy explicitly as anyons nontrivially linking themselves in the bulk, as shown in [274, *Gu:2016hoy].
@@
This point of view of understanding quantum chaos and OTOC from the exchange algebra was emphasized in the context of 2D CFT in [57, 276, 277]. In the context of holography, this is the dual of the bulk exchange algebra found in [278, *Kiem:1995iy], responsible for the exponential blueshift coming from the Dray and ’t Hooft shockwave effect[280]. As noted in [57], the goal of conformal bootstrap is to ensure that correlation functions remain single-valued in Euclidean signature. In other words, quantum statistics are invisible or “confined” in Euclidean, non-chiral, full CFTs. However, after we solve the bootstrap equations which are topological in nature as we explained, and manage to hide the quantum statistics, these topological data will get “deconfined” in Lorentzian signature, controlling the Lorentzian dynamics, causality(commutator), and information scrambling of the theory, which gives “Chaos from Order”.
@@ 7 Discussions
As discussed above, the bulk dual of the CFT is represented by a sum over quantum geometries, providing an exact description at finite $N$. In the large $N$ limit, certain quantum geometries emerge as dominant saddles in this sum, allowing for a semi-classical, low-energy geometrical description. To better understand this emergence, recent studies have proposed algebraic quantum field theory (AQFT), specifically von Neumann algebras, as a suitable framework, noting the appearance of emergent Type III von Neumann algebra structures in the large $N$ limit[282, *Leutheusser:2021frk].
A natural question arises: how are these two distinct algebraic approaches for field theories and emergence of geometry related, and how are categorical structures, such as the $6j$ symbols, encoded within von Neumann algebras? In 2D CFT, there appears to be a dictionary connecting these approaches. AQFT in 2D is rigorously constructed using “conformal nets”, and its connection to the functorial formalism is discussed in [284]. A key element in this framework is the role of superselection sectors and “subfactor theory” [285, *Longo:1989tt] [^51] We thank Keiichiro Furuya for discussions on this topic and collaboration on related project.. This approach might provide a new understanding of the Type II algebra structures observed in the context of higher-dimensional gravity [287], which plays a role in black holes and cosmology [288, *Chandrasekaran:2022cip, 290]. Applied to 3D gravity, this Type II algebra is analogous to the JT gravity case studied in [291, *Kolchmeyer:2023gwa]. In the current situation, we even achieve its UV completion, since we start from a 2D CFT. We will report further results on these ideas in future work.
@@
In the case of irrational CFTs, an additional layer of infinity appears due to the infinite number of primary fields in |$\ket{\Psi}$. This infinity drives the large $N$ limit of a CFT and enables chaotic behavior, in contrast to rational CFTs. It mirrors the emergence of Type III factors in the large $N$ limit as discussed above [282, *Leutheusser:2021frk]. For the SymTFT description, this implies that an infinite number of anyons are required to describe the bulk theory of gravity. Consequently, for irrational CFTs, we do not have a fusion category, which by definition requires finitely many objects. This fact also limits our Turaev-Viro construction for $U_{q}(SL(2,\mathbb{R}))$, as it lacks the 4-1 move invariance. However, when we consider these three-dimensional manifolds with boundaries, combined with boundary conformal blocks, the construction does make sense and yields a convergent result consistent with the Virasoro/Teichmuller TQFT, at least for the diagonal Liouville CFT, as explicitly shown in [1, 27]. This suggests the existence of a novel mathematical structure, essential for the study of quantum gravity.
@@
Our tensor network formalism depends significantly on the unique properties of 3D gravity and 2D CFT, particularly its topological nature. Extending this line of reasoning may provide insight into constructing a non-perturbative bulk dual in higher dimensions. However, since gravitons are dynamical in higher dimensions, we want to explore several possibilities to address this tension. One possibility, as noted in [39], is that the bulk theory may lose its topological character due to the irrational nature of the symmetry. A second possibility is that the bulk topological invariance could be spontaneously broken, even if it initially exhibits a topological structure. Alternatively, it is possible that the bulk theory remains genuinely topological in the non-perturbative regime. These two possibilities might be related if we expand around fixed backgrounds.
@@
The volume/action of the bulk theory is proposed to be connected to the notion of complexity in the CFT [297, *Stanford:2014jda, *Brown:2015bva]. There is a large volume of work exploring this connection, and it has been shown that perhaps more generally there are an infinite class of bulk quantities that could be understood as complexity[300, *Belin:2022xmt, *Myers:2024vve]. However, precise definitions of complexity on both sides have been tricky to identify. Our discrete formulation of the path-integral could be used to define notions of complexity in field theories that can be directly translated between the bulk and the boundary. For example, since each 6j symbol for the Virasoro lines can be interpreted as the volume of hyperbolic tetrahedra, we can assign definite complexity to individual “gate” corresponding to the 6j symbols that builds up the path-integral. Our $U_{q}(SL(2,\mathbb{R}))$ $6j$ symbols seem to be particularly promising, since Liouville theory has be proposed to capture path integral complexity for 2D CFTs[303, *Caputa:2017yrh].
@@ Acknowledgement
We thank Nathan Benjamin, Christian Ferko, Keiichiro Furuya, Hao Geng, Sarah Harrison, Daniel Jafferis, Hong Liu, Conghuan Luo, Juan Maldacena, John McGreevy, Shinsei Ryu, Sahand Seifnashri, Shu-Heng Shao, Herman Verlinde and Frank Verstraete for helpful discussions. We thank Wei Bu, Wan Zhen Chua, Christian Ferko, Hao Geng, Ziming Ji, Chen-Te Ma, Yixu Wang, Zixia Wei and Chen Yang for comments and discussions on the draft. We thank Lin Chen, Gong Cheng, Keiichiro Furuya, Hao Geng, Zheng-Cheng Gu, Bing-Xin Lao for collaboration on related projects. We thank Zhenhao Zhou for plotting the 3D diagrams in this paper. YJ thanks Juan Maldacena for the invitation and hospitality during a visit to the Institute for Advanced Study, where the idea of this paper originated from discussions with people there. YJ and NB acknowledge the support by the U.S Department of Energy ASCR EXPRESS grant, Novel Quantum Algorithms from Fast Classical Transforms, and Northeastern University. LYH acknowledges the support of NSFC (Grant No. 11922502, 11875111). ZL is supported by NSF grant PHY-2110463.
@@ A.1 Turaev-Viro TQFT from Chern-Simons theory
By considering the dual graph corresponding to the network of Wilson lines, and recursively using the relation in Fig. 33, it recovers the path integral with the insertion of $\omega$ loops, which serves only to project every loop back to vanishing flux[307, *2011NJPh...13f5001B]. The resultant network is called the “Chain-Mail” [305]. The corresponding expectation value of closed Wilson loops is denoted as $C(\mathcal{M},D)$. Roberts invariants can be defined as:
@@ A.2 Turaev-Viro TQFT from fusion category
Generically, it is possible to endow a fusion category with more structures. An important structure is the braided structure that exchanges two objects. This structure is encoded in the so-called $R$-matrix, with components $R^{c}_{ab}$ which are phases, that connects $a\otimes b$ to $b\otimes a$. This is illustrated in Fig. 37. Note that there is a distinction between an “over-crossing” and “under-crossing”. The factor that appears in an “under-crossing” is related to the “over-crossing” by a complex conjugate of $R_{ab}^{c}$. The $R$-matrix satisfies some consistency conditions, so that the action of braiding is consistent with the associativity relations from fusion. These consistency conditions are given by the Hexagon relations[52, *Moore:1989vd].
@@ Appendix B Tensor network state sum representation of 2D CFT path integrals
For the computation of the total CFT partition $\mathcal{Z}_{\text{CFT}}$ function, the basic building block is the CFT path integral on the chipped triangles, as shown in Fig. 9555555We will focus on a 2D CFT defined on a flat background $\delta_{ij}$ for simplicity. However, the formalism can be easily generalized to general fixed curved backgrounds $h_{ij}$. For many 2D manifolds that do not admit a global flat metric, such a generalization must be applied.. These represent open pairs of pants, in contrast to the more conventional closed pairs of pants commonly found in the CFT literature. The Hilbert spaces in the computation are associated with the edges of the triangles, which are lines between the two endpoints where boundary conditions must be specified.
@@
For diagonal CFTs, the label set of conformal boundary conditions coincides with the set of primary fields (this is referred to as the Cardy case [103] in the literature), and the rescaled BCFT OPE coefficients are directly proportional to the quantum $6j$ symbols, as first discovered by Runkel [99, *Runkel:1999dz], as
@@
where $d_{\alpha_{i}}$ is the quantum dimension associated with $\alpha_{i}$, as introduced above. For theories with a continuous spectrum, such as the Liouville theory, this corresponds to the Plancherel measure [1, 156, *Ponsot:2000mt]. In the diagonal case, the $6j$ symbol enjoys the tetrahedron symmetry, and we associate it with a tetrahedron, which serves as the fundamental building block of the Turaev-Viro TQFT, as reviewed above.
@@ Appendix C Tensor network SymQRG
The orthogonality condition of the (generalized) $6j$ symbols is[28, *Fuchs:2004xi, 101],
@@
The next step is to express the wavefunctions as contractions of local tensors defined on the triangles as in Fig. 45, using singular value decomposition [109, *PhysRevLett.115.180405, 25], as follows:
@@
In the context of 3D gravity, the idea of seeing the emergence of gravity by treating closed OPE coefficients in the pair-of-pants decomposition as random tensors was introduced in [316, *Verlindeunpublished], and later adapted into a tensor network framework in [318]. For example, a simple MPS tensor network, constructed with random closed OPE coefficients in a 2D CFT as tensors, can be interpreted as discretizing the radial direction in the bulk dual in the semi-classical approximation, and approximately reproduces the location of the horizon[316, *Verlindeunpublished, 318]. In our framework, which recovers the full CFT path integral, it is also natural to consider an ensemble average of the structure coefficients for boundary-changing operators.
@@
The ensemble average of the structure coefficients for the boundary-changing operators in the large central charge limit has been studied in [187, *Kusuki:2022wns, 189]. Applied to our formalism, this results in a random tensor network that represents the ensemble-averaged CFT path integral as outlined in [1], and will appear soon in [191]. A crucial distinction in our tensor network construction is worth emphasizing: We used the open structure coefficients of the CFT, rather than the closed ones, to build the CFT tensor network. From our perspective, the open structure coefficients are more directly connected to the emergence of bulk geometry and allow for discretizing both the CFT and AdS spaces arbitrarily, not only along the radial direction or having to come with external operator insertions.